# Deep Reinforcement Learning for Dynamic Algorithm Selection: A Proof-of-Principle Study on Differential Evolution

### 06 Mar 2024 v2

@@ Abstract
Index Terms:
Algorithm selection, deep reinforcement learning, meta-black-box optimization, black-box optimization, differential evolution.
I Introduction
@@
The rest of this paper is organized as follows: Section II presents the preliminaries on DRL and DE. Section III introduces the detail of RL-DAS. Section IV is the proof-of-principle study of applying RL-DAS on advanced DE algorithms. The experimental results are depicted in Section V with detailed analysis, followed by a conclusion in Section VI.
II Preliminaries and Related Works
In this section, we briefly introduce the Markov Decision Process and the advanced variants of DE. Table I summarizes the abbreviation indices for easy reference.
TABLE I: The abbreviation indices.
@@ II Preliminaries and Related Works
II-A Markov Decision Process and Policy Gradient
@@ II-B3 RL-Assisted DE
III Methodology
III-A Dynamic Algorithm Selection
@@ III-B RL-DAS Overview
In the following of this section, we introduce the design of MDP components including the state, action and reward in Section III-C. The neural network design and the training process will be presented in Section III-D. Last, we introduce the aforementioned algorithm context memory mechanism to support the warm-start of different algorithms in Section III-E.
@@ III-D3 Critic
III-E Algorithm Context Memory
@@
IV RL-DAS on DE: A Proof-of-Principle Study
The design of RL-DAS can be applied to various EC algorithms, including GA, PSO, DE, and their mixtures, each with corresponding algorithm pools. However, in this paper, for the sake of convenience in the comparison and discussion, we focus on presenting a proof-of-principle study using representative DE algorithms. As mentioned in Section II-B, DE is a high-profile field in EC. Although exhibits powerful performance in continuous optimization, different DE variants generally exhibit different performances on different problem instances. We realize the above RL-DAS framework to dynamically schedule three DE algorithms along the optimization process, which is detailed in this section.
IV-A Candidate Algorithm Pool
We adopt three advanced DE algorithms from the CEC 2021 competition [14], namely, JDE21 [21], MadDE [25] and NL-SHADE-RSP [23]. These three algorithms have been introduced in Section II-B, which have different strengths and weaknesses in terms of balancing the exploration and exploitation of the population. According to our experiments (which will be discussion in Section 3 of our supplementary document), these three algorithms present different search characteristics and exhibit complementary performance in different optimization stages of different problem instances, which leaves room for our RL-DAS agent to learn a desired collaboration between them.
IV-B Schedule Interval
@@
According to the RL practice, the common length of an RL trajectory ranges from $T_{\text{min}}=10$ to $T_{\text{max}}=200$ [45]. To this end, considering the optimization length in EC, which is usually controlled by the $MaxFEs$, the suggested schedule interval of RL-DAS is $\varDelta\in\left[\frac{MaxFEs}{T_{\max}},\frac{MaxFEs}{T_{\min}}\right]$. In essence, at every $\varDelta$ FEs, the framework attempts to insert a transition (algorithm switching) step of RL. In practical execution, the framework will wait until the current generation of the DE algorithm completed before starting the transition step. As a result, the real scheduling interval, $\tilde{\varDelta}$, may be equal to or slightly larger than the planned $\varDelta$. In Section V-E, we will conduct a grid experiment to investigate interval lengths that support good performance with acceptable training efficiency.
IV-C Algorithm Restoration
@@ V Experiment
  • RQ1: How does the performance of our RL-DAS compare with its backbone candidate algorithms, an optimal estimate of traditional AS, and other advanced DE variants (including adaptive, ensemble, and RL-assisted ones)? See Section V-C.
  • RQ2: How well does RL-DAS generalize to unseen problem instances? See Section V-D.
  • RQ3: What are the suggested configurations for RL-DAS? See Section V-E.
V-A Benchmark Generation
@@ V-C Performance Comparisons
TABLE II: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 10D problems.
@@
TABLE III: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 20D problems.
@@
V-C1 Comparison with Backbones
Table II and Table III show the performance of algorithms on 10D and 20D problems respectively, where the upper value is the descent which describes the average descent of costs in percentage (normalized by the best costs in the initial populations, which are consistent among algorithms with random number seeds), and the lower value is FEs describes the average consumed function evaluations to reach the termination criterion. The ‘$+$’, ‘--’ and $\approx$’ indicate the statistical test results that the performance of competing methods is better, worse and no significant difference with RL-DAS respectively according to the Wilcoxon rank-sum test at the 0.05 significance level. The last columns of the two tables summarize the statistical results where the first value counts the cases when RL-DAS significantly outperforms the competitor (the two tests contain at least one ‘--’ while the test on descent is not ‘$+$’); the second value counts the cases that there is no significant difference between RL-DAS and the competitor (both tests are $\approx$’); and the third value counts the cases otherwise. From the results in these tables, we can conclude that:
@@ V-C1 Comparison with Backbones
  • For the 20D problems shown in Table III, RL-DAS still outperforms the others. The increased problem space makes the optimization more difficult, and our relative performance over the competitors is even better than that on 10D problems. The results indicate that dimensions do not affect the RL-DAS much, the conclusions on 10D problems are still applicable to the 20D benchmark.
@@
V-C2 Comparison between DAS and AS
The DAS and backbones parts of Table II and Table III verify that RL-DAS outperforms the compared single algorithm. Then a question comes out: can RL-DAS defeat other methods that combine these single algorithms, for example, by the Algorithm Selection (AS)? To figure out the answer, we introduce a competitor that takes the best result among the three backbone algorithms as its performance, namely $*{}^{*}$, to simulate the performance upper bound of AS methods. For each problem instance $i$, the result of $*{}^{*}$ is calculated by:
@@ V-C2 Comparison between DAS and AS
where JDE21($i$) indicates the optimization result of JDE21 on instance $i$, and so do the items for MadDE and NL-SHADE-RSP. The performance comparison is shown in the AS parts of Table II and Table III.
It can be seen that RL-DAS outperforms $*{}^{*}$ in most of the problems. AS methods employ a single candidate algorithm for each problem instance, therefore the upper bound of its performance can not exceed the best performance among all candidate algorithms, let alone that practical AS methods usually require external evaluation resources to select an algorithm [6] which will inevitably impact the performance. Differently, the DAS methods select an algorithm at each interval which has an opportunity to present a complementary performance and breaks through the single best candidate algorithm. The outstanding performance of Rand-DAS in Table II and Table III prove its potential. The employment of an DRL agent ensures wise selection thus RL-DAS defeats $*{}^{*}$ significantly.
V-C3 Comparison with Other Advanced DE Variants
The last part in Table II and Table III present the comparison between our RL-DAS and other advanced DE. For the sake of fairness, the three RL-assisted methods are trained on the same training set as RL-DAS and update their models by equal steps.
@@ V-C3 Comparison with Other Advanced DE Variants
TABLE IV: Computational cost comparison.
@@
V-C4 Runtime Comparison
Table IV presents the average computational time of different algorithms on 10D C2 (Schwefel) testing set problems. The upper row represents traditional non-learning algorithms, while the lower row presents the RL-assisted algorithms including our RL-DAS. It is evident that the computational costs of most DE algorithms are comparable, with MadDE being relatively slower due to its use of more operators and parameter adaptive mechanisms. RL-assisted algorithms, in general, require more time for network inference. Among them, DEDDQN takes the longest time, attributed to its costly and complex 99-dimensional state representation and one-by-one operator selection for each individual. RL-DAS, on the other hand, completes optimization in 1.434 seconds for 200,000 FEs. While it requires more time for network inference and algorithm switching compared to JDE21, MadDE, and NL-SHADE-RSP, the additional computational cost is generally acceptable regarding the state-of-the-art optimization accuracy of RL-DAS.
@@ V-C4 Runtime Comparison
TABLE V: Transfer results from one class of problems to others.
@@
TABLE VI: Transfer results under different problem class partitions.
@@
V-D Zero-shot Knowledge Transfer
In this section, we conduct two types of zero-shot knowledge transfer experiments. First, we choose models from the three problem classes which have significant performance differences as reported in Section V-C: the unimodal problem C1 (Bent Cigar), the basic multimodal problem C2 (Schwefel) and the mix-class problems C11, and then zero-shot transfer them to the other CEC problems. The results are shown in Table V. It can be seen that RL-DAS successfully transfers its learned knowledge to unseen problems, which allows the models from C1, C2 and C11 achieve close results to the original RL-DAS.
In the second experiment, we establish different transfer settings, denoted as TS1-TS3, by partitioning the ten classes of problems into training and testing sets with varying proportions: 2 : 8 (F1-F2 for training, F3-F10 for testing), 5 : 5 (F1-F5 for training, F6-F10 for testing), and 8 : 2 (F1-F8 for training, F9-F10 for testing), respectively. The training and testing sets, with instances mixed from different problem classes, are generated in the same way as C11. The results are shown in Table VI, where the three backbone algorithms serve as references. RL-DAS outperforms the baselines on most of transfer settings in Table VI. It demonstrates promising generalization ability, transferring the knowledge learned from a small training problem set to a much larger and more complex problem set.
@@ V-D Zero-shot Knowledge Transfer
V-E Ablation Study and Hyperparameter Research
V-E1 Ablation Study
@@
Table VII shows the results. As can be observed, the significant role of the two features is highlighted by the poor performance if they are removed. For the ablation of the context memory, the results also demonstrate a significant degradation in performance. Without historical memory restoration, algorithms are compelled to optimize the population from scratch in each period, hindering comprehensive algorithm cooperation
TABLE VII: The ablation study performance.
@@ V-E2 Reward Design
TABLE VIII: The performance under different reward schemes.
@@
The comparison results for 10D problems are presented in Table VIII. Our reward explicitly considers both cost and speed, offering valuable guidance to the agent. It is evident from the table that our reward outperforms others on the majority of problems, particularly in terms of optimization speed. In contrast, alternative reward functions such as $r1$ offer no differentiation on FE consumption, resulting in identical rewards for fast or slow convergence. Consequently, the agent lacks motivation to expedite optimization, leading to relatively poor performance.
V-E3 Schedule Interval
@@ Appendix B Experiment Settings
TABLE IX: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 10D problems.
@@
Appendix C In-Depth Analysis of DAS
Figure 5: The descent curve (a) and selection distribution (b) on C2.
TABLE X: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 20D problems.
@@ Appendix C In-Depth Analysis of DAS
TABLE XI: Transfer results from one class of problems to others.
@@
TABLE XII: Transfer results under different problem class partitions.
@@
TABLE XIII: The ablation study performance.
@@
TABLE XIV: The optimization performance under different reward schemes.
@@ Appendix D Absolute Cost Values
In the main body of our paper, we report the cost decent in percentage of different algorithms from the random initialization. This section additionally presents the detailed absolute results of the algorithms, where the mean and standard deviations are presented. Specifically, Table IX and Table X supplement the results for major performance comparisons of RL-DAS and the competitors. Table XI and Table XII shows the detailed results of the zero-shot transfer experiment. Table XIII is the results for the ablation study on framework components, while Table XIV delves into the results of the reward study.

### 07 Mar 2024 v3

@@ Abstract
Keywords– Algorithm selection, deep reinforcement learning, meta-black-box optimization, black-box optimization, differential evolution
1 Introduction
@@
The rest of this paper is organized as follows: Section 2 presents the preliminaries on DRL and DE. Section 3 introduces the detail of RL-DAS. Section 4 is the proof-of-principle study of applying RL-DAS on advanced DE algorithms. The experimental results are depicted in Section 5 with detailed analysis, followed by a conclusion in Section 6.
2 Preliminaries and Related Works
In this section, we briefly introduce the Markov Decision Process and the advanced variants of DE. Table 1 summarizes the abbreviation indices for easy reference.
Table 1: The abbreviation indices.
@@ 2 Preliminaries and Related Works
2.1 Markov Decision Process and Policy Gradient
@@ 2.2.3 RL-Assisted DE
3 Methodology
3.1 Dynamic Algorithm Selection
@@ 3.2 RL-DAS Overview
In the following of this section, we introduce the design of MDP components including the state, action and reward in Section 3.3. The neural network design and the training process will be presented in Section 3.4. Last, we introduce the aforementioned algorithm context memory mechanism to support the warm-start of different algorithms in Section 3.5.
@@ 3.4.3 Critic
3.5 Algorithm Context Memory
@@
4 RL-DAS on DE: A Proof-of-Principle Study
The design of RL-DAS can be applied to various EC algorithms, including GA, PSO, DE, and their mixtures, each with corresponding algorithm pools. However, in this paper, for the sake of convenience in the comparison and discussion, we focus on presenting a proof-of-principle study using representative DE algorithms. As mentioned in Section 2.2, DE is a high-profile field in EC. Although exhibits powerful performance in continuous optimization, different DE variants generally exhibit different performances on different problem instances. We realize the above RL-DAS framework to dynamically schedule three DE algorithms along the optimization process, which is detailed in this section.
4.1 Candidate Algorithm Pool
We adopt three advanced DE algorithms from the CEC 2021 competition [14], namely, JDE21 [21], MadDE [25] and NL-SHADE-RSP [23]. These three algorithms have been introduced in Section 2.2, which have different strengths and weaknesses in terms of balancing the exploration and exploitation of the population. According to our experiments (which will be discussion in Section 3 of our supplementary document), these three algorithms present different search characteristics and exhibit complementary performance in different optimization stages of different problem instances, which leaves room for our RL-DAS agent to learn a desired collaboration between them.
4.2 Schedule Interval
@@
According to the RL practice, the common length of an RL trajectory ranges from $T_{\text{min}}=10$ to $T_{\text{max}}=200$ [45]. To this end, considering the optimization length in EC, which is usually controlled by the $MaxFEs$, the suggested schedule interval of RL-DAS is $\varDelta\in\left[\frac{MaxFEs}{T_{\max}},\frac{MaxFEs}{T_{\min}}\right]$. In essence, at every $\varDelta$ FEs, the framework attempts to insert a transition (algorithm switching) step of RL. In practical execution, the framework will wait until the current generation of the DE algorithm completed before starting the transition step. As a result, the real scheduling interval, $\tilde{\varDelta}$, may be equal to or slightly larger than the planned $\varDelta$. In Section 5.5, we will conduct a grid experiment to investigate interval lengths that support good performance with acceptable training efficiency.
4.3 Algorithm Restoration
@@ 5 Experiment
  • RQ1: How does the performance of our RL-DAS compare with its backbone candidate algorithms, an optimal estimate of traditional AS, and other advanced DE variants (including adaptive, ensemble, and RL-assisted ones)? See Section 5.3.
  • RQ2: How well does RL-DAS generalize to unseen problem instances? See Section 5.4.
  • RQ3: What are the suggested configurations for RL-DAS? See Section 5.5.
5.1 Benchmark Generation
@@ 5.3 Performance Comparisons
Table 2: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 10D problems.
@@
Table 3: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 20D problems.
@@
5.3.1 Comparison with Backbones
Table 2 and Table 3 show the performance of algorithms on 10D and 20D problems respectively, where the upper value is the descent which describes the average descent of costs in percentage (normalized by the best costs in the initial populations, which are consistent among algorithms with random number seeds), and the lower value is FEs describes the average consumed function evaluations to reach the termination criterion. The ‘$+$’, ‘--’ and $\approx$’ indicate the statistical test results that the performance of competing methods is better, worse and no significant difference with RL-DAS respectively according to the Wilcoxon rank-sum test at the 0.05 significance level. The last columns of the two tables summarize the statistical results where the first value counts the cases when RL-DAS significantly outperforms the competitor (the two tests contain at least one ‘--’ while the test on descent is not ‘$+$’); the second value counts the cases that there is no significant difference between RL-DAS and the competitor (both tests are $\approx$’); and the third value counts the cases otherwise. From the results in these tables, we can conclude that:
@@ 5.3.1 Comparison with Backbones
  • For the 20D problems shown in Table 3, RL-DAS still outperforms the others. The increased problem space makes the optimization more difficult, and our relative performance over the competitors is even better than that on 10D problems. The results indicate that dimensions do not affect the RL-DAS much, the conclusions on 10D problems are still applicable to the 20D benchmark.
@@
5.3.2 Comparison between DAS and AS
The DAS and backbones parts of Table 2 and Table 3 verify that RL-DAS outperforms the compared single algorithm. Then a question comes out: can RL-DAS defeat other methods that combine these single algorithms, for example, by the Algorithm Selection (AS)? To figure out the answer, we introduce a competitor that takes the best result among the three backbone algorithms as its performance, namely $*{}^{*}$, to simulate the performance upper bound of AS methods. For each problem instance $i$, the result of $*{}^{*}$ is calculated by:
@@ 5.3.2 Comparison between DAS and AS
where JDE21($i$) indicates the optimization result of JDE21 on instance $i$, and so do the items for MadDE and NL-SHADE-RSP. The performance comparison is shown in the AS parts of Table 2 and Table 3.
It can be seen that RL-DAS outperforms $*{}^{*}$ in most of the problems. AS methods employ a single candidate algorithm for each problem instance, therefore the upper bound of its performance can not exceed the best performance among all candidate algorithms, let alone that practical AS methods usually require external evaluation resources to select an algorithm [6] which will inevitably impact the performance. Differently, the DAS methods select an algorithm at each interval which has an opportunity to present a complementary performance and breaks through the single best candidate algorithm. The outstanding performance of Rand-DAS in Table 2 and Table 3 prove its potential. The employment of an DRL agent ensures wise selection thus RL-DAS defeats $*{}^{*}$ significantly.
5.3.3 Comparison with Other Advanced DE Variants
The last part in Table 2 and Table 3 present the comparison between our RL-DAS and other advanced DE. For the sake of fairness, the three RL-assisted methods are trained on the same training set as RL-DAS and update their models by equal steps.
@@ 5.3.3 Comparison with Other Advanced DE Variants
Table 4: Computational cost comparison.
@@
5.3.4 Runtime Comparison
Table 4 presents the average computational time of different algorithms on 10D C2 (Schwefel) testing set problems. The upper row represents traditional non-learning algorithms, while the lower row presents the RL-assisted algorithms including our RL-DAS. It is evident that the computational costs of most DE algorithms are comparable, with MadDE being relatively slower due to its use of more operators and parameter adaptive mechanisms. RL-assisted algorithms, in general, require more time for network inference. Among them, DEDDQN takes the longest time, attributed to its costly and complex 99-dimensional state representation and one-by-one operator selection for each individual. RL-DAS, on the other hand, completes optimization in 1.434 seconds for 200,000 FEs. While it requires more time for network inference and algorithm switching compared to JDE21, MadDE, and NL-SHADE-RSP, the additional computational cost is generally acceptable regarding the state-of-the-art optimization accuracy of RL-DAS.
@@ 5.3.4 Runtime Comparison
Table 5: Transfer results from one class of problems to others.
@@
Table 6: Transfer results under different problem class partitions.
@@
5.4 Zero-shot Knowledge Transfer
In this section, we conduct two types of zero-shot knowledge transfer experiments. First, we choose models from the three problem classes which have significant performance differences as reported in Section 5.3: the unimodal problem C1 (Bent Cigar), the basic multimodal problem C2 (Schwefel) and the mix-class problems C11, and then zero-shot transfer them to the other CEC problems. The results are shown in Table 5. It can be seen that RL-DAS successfully transfers its learned knowledge to unseen problems, which allows the models from C1, C2 and C11 achieve close results to the original RL-DAS.
In the second experiment, we establish different transfer settings, denoted as TS1-TS3, by partitioning the ten classes of problems into training and testing sets with varying proportions: 2 : 8 (F1-F2 for training, F3-F10 for testing), 5 : 5 (F1-F5 for training, F6-F10 for testing), and 8 : 2 (F1-F8 for training, F9-F10 for testing), respectively. The training and testing sets, with instances mixed from different problem classes, are generated in the same way as C11. The results are shown in Table 6, where the three backbone algorithms serve as references. RL-DAS outperforms the baselines on most of transfer settings in Table 6. It demonstrates promising generalization ability, transferring the knowledge learned from a small training problem set to a much larger and more complex problem set.
@@ 5.4 Zero-shot Knowledge Transfer
5.5 Ablation Study and Hyperparameter Research
5.5.1 Ablation Study
@@
Table 7 shows the results. As can be observed, the significant role of the two features is highlighted by the poor performance if they are removed. For the ablation of the context memory, the results also demonstrate a significant degradation in performance. Without historical memory restoration, algorithms are compelled to optimize the population from scratch in each period, hindering comprehensive algorithm cooperation
Table 7: The ablation study performance.
@@ 5.5.2 Reward Design
Table 8: The performance under different reward schemes.
@@
The comparison results for 10D problems are presented in Table 8. Our reward explicitly considers both cost and speed, offering valuable guidance to the agent. It is evident from the table that our reward outperforms others on the majority of problems, particularly in terms of optimization speed. In contrast, alternative reward functions such as $r1$ offer no differentiation on FE consumption, resulting in identical rewards for fast or slow convergence. Consequently, the agent lacks motivation to expedite optimization, leading to relatively poor performance.
5.5.3 Schedule Interval
@@ Appendix B Experiment Settings
Table 9: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 10D problems.
@@
Table 10: Comparing RL-DAS with comparison algorithms on Shifted & Rotated 20D problems.
@@
Appendix C In-Depth Analysis of DAS
Figure 5: The descent curve (a) and selection distribution (b) on C2.
Table 11: Transfer results from one class of problems to others.
@@ Appendix C In-Depth Analysis of DAS
Table 12: Transfer results under different problem class partitions.
@@
Table 13: The ablation study performance.
@@
Table 14: The optimization performance under different reward schemes.
@@ Appendix D Absolute Cost Values
In the main body of our paper, we report the cost decent in percentage of different algorithms from the random initialization. This section additionally presents the detailed absolute results of the algorithms, where the mean and standard deviations are presented. Specifically, Table 9 and Table 10 supplement the results for major performance comparisons of RL-DAS and the competitors. Table 11 and Table 12 shows the detailed results of the zero-shot transfer experiment. Table 13 is the results for the ablation study on framework components, while Table 14 delves into the results of the reward study.